Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 40 miles per hour and train B is traveling at 48 miles per hour. Train A passes a station ar 1:25 a.m. If train B passes the same station at 1:40 a.m., at what time will train B catch up to train A?
Answer by mananth(16946)

Train A speed = 40 mph
Train B 48 mph
Train A is ahead of train B at the station by 15 minutes.( 1/4 hour)
By the time train B reaches the station train A has traveled 40 * 1/4 miles
=10 miles to a point say P.
Let the distance from point P to the point where train B overtakes train A be x
.
Train B travels 10 + x miles in the same time the train A travels x miles
time taken by train A = x/40 hours
time taken by train B = x+10 / 48
.
x/40 = x+10 /48
48x= 40(x+10)
48x=40x +400
8x= 400
x= 50
time taken by train A t travel 50 miles = 50/40
=5/4 hours= 75 minutes
This is the same time taken by train B to cross train A
Add 75 minutes to the time when Train B was at the station. You know what to do
That is the time train B catches up